OpenStudy (refrac532):

The graph of a cosine function has amplitude = 3, period = 90 degrees, and a vertical translation 2 units up. Write an equation for the function. Then sketch the graph.

OpenStudy (mww):

If a cosine function has amplitude of A, and a period of T, shifted up K units, the equation is y = A cos(bx) + K Now period is given by T = 2pi/b so b = 2pi/T So sub in your amplitude, b and your K into the formula.

OpenStudy (mww):

as I see you are using degree measure rather than radians, you can use T = 360/b instead so b = 360/T